Calculate the area of a rectangle whose length and breadths are given as 3x2y m and 5xy2 m respectively.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,

Length = 3x^2 y m

Breadth = 5x y^2m

Area of rectangle = Length × Breadth

= (3x^2^y × 5x y^2) = (3 × 5) × x^2 y × x y^2= 15x^3y^3m^2
thanks